OpenStudy (ttp):

If a leaky faucet wastes 50 gallons in five weeks, about how many gallons does it waste over a year (52 weeks)? ten gallons two hundred sixty gallons four hundred seventy gallons five hundred twenty gallons

OpenStudy (anonymous):

D

OpenStudy (texaschic101):

proportion time.. 50/5 = x/52 (50 gal to 5 weeks = x gal to 52 weeks) cross multiply (5)(x) = (52)(50) 5x = 2600 x = 2600/5 x = 520

OpenStudy (texaschic101):

helpout is correct....answer is D

OpenStudy (anonymous):

Sorry i did the math in my head i should done it on here

OpenStudy (ttp):

Thanks @helpout3 and @texaschic101 :)

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Yeah no problem anytime if i can actually help